OpenStudy (anonymous):

formulate an equation and solve. four times the sum of three times a number increased by two is equal to three times the difference of two times the number decreased by eight. what is the number?

OpenStudy (anonymous):

hello an example 5 times the sum of 2 times a number increased by 10 5 x ( 2 n + 10 )

OpenStudy (anonymous):

4(3x+2)=3(2x-8) 12x+8=6x-24 6x=-32 x=-32/6 or -5 1/3

OpenStudy (anonymous):

excellent

OpenStudy (anonymous):

It was a pleasure to help you